4. Search with imperfect information After graduating, you start work as a management consultant. You are paid $150 per hour one morning before work, you decide to buy a new car. $35,500 to $36,500 with the You know the exact model you w nd you know that in your area the price ranges from average price ant, a you can expect to get being $36,000. You can choose among hundreds of dealers, but you dont know which dealer will give you the best price. Time is literally money, since every hour you spend searching is an hour you dont get paid. Each visit to a dealer takes an hour Your expected marginal benefit of another search is the difference between the current dealers offer and the average price. to asks $36,300 for the c Should you The first dealer pt the price or keep searching? (Keep in mind that each visit to a dealer takes an you go hour O Keep searching O Accept the price. to asks $36,100. Do you think you should accept this price or keep searching? Suppose you kept searching, a nd the next dealer you go O Keep searching O Accept the price. Suppose you kept searching, a nd the next dealer to asks $36,130. You could return to the last dealer (who offered you a price of $36,100) but you go that would take another hour What should you do? O Return to the last dealer and pay $36,100. O Get a new price from yet another dealer O Accept the price of $36,130 which of the following best explains how it can be profitable for different dealers to offer different prices O Some dealers have lower costs than others. O Some dealers have higher costs than others It he higher-priced dealers will lose money in the long run isnt-t O Because different people have different opportunity costs of their time, even high-priced dealers can sometimes make sales

 
Time is literally money, since every hour you spend searching is an hour you don't get paid. Each visit to a dealer takes an hour. Your expected marginal benefit of another search is the difference between the current dealer's offer and the average price.
 
The first dealer you go to asks $36, 300, for the car. Should you accept the price or keep searching? (Keep in mind that each visit to a dealer takes an hour, ) Keep searching. Accept the price.
 
Suppose you kept searching, and the next dealer you go to asks $36, 100. Do you think you should accept this price or keep searching? Keep searching. Accept the price.
 
Suppose you kept searching, and the next dealer you go to asks $36, 130. You could return to the last dealer (who offered you a price of $36, 100) but that would take another hour. What should you do? Return to the last dealer and pay $36, 100. Get a new price from yet another dealer. Accept the price of $36, 130.
 
Which of the following best explains how it can be profitable for different dealers to offer different prices? Some dealers have lower costs than others. Some dealers have higher costs than others. It isn't-the higher-priced dealers will lose money in the long run. Because different people have different opportunity costs of their time, even high-priced dealers can sometimes make sales.
Answer #1

We know the Average price of the car is $36,000.

Each dealer takes 1 hour that means going to each dealer costs me $150 as opportunity cost. ( Because instead of finding the suitable dealer, I could work and earn $150 per hour)

Case 1:

Dealer asks for $36,300. Price is above the average price and Including my opportunity costs, it will cost me $36,300 + 150 = $36,450

I would like to go to another dealer i.e. KEEP SEARCHING.

Case 2:

Dealer asks for $36,100. Price is above the average price and Including my opportunity costs, it will cost me $36,100 + 150 + 150 = $36,400

It's better than Case 1. It would like to accept it.

Because, even if next dealer asks me the average price that is 36,000, it would cost me ( 36,000 + 150 *3)36,450. So, the better option is to accept it.

Case 3:

Suppose kept searching, next dealer asks for $36,130. I have an option even to get back to the previous seller, so what would I like to do.

I will accept this price.

Because, third dealer's price is costing me $36,130 + 150*3 = $36,580

AND returning to the second seller will cost me another 1 hour, so second dealer's price is costing me now $36,100 + 150*4 = $36,700

And searching for another dealer even if he asks for $36,000 , it will cost me $36,000 + 150 * 4 = $36,600

So, better option among these 3 is to accept the price of the third dealer.

*** it can be profitable for different dealers to offer different prices because different people have different opportunity costs of the time, even high priced dealers can sometimes make sales like in case 3. ***